OpenStudy (anonymous):

The cost of cementing a walk 8 feet wide and 24 feet long at $14.40 a square yard is

OpenStudy (anonymous):

i'm not sure if it is $921.6 because 8*24/3=64 and 64(14.4)= 921.6 OR 8/3(8) which is 64/3 yd(14.4) = $307.20. Please help!

OpenStudy (triciaal):

so we need to find the area of a rectangle 8ft * 24ft the cost is per square yard so we need to convert the area found in square ft to square yard 3 ft = 1 yd so 1 square yd = 9 square ft area in sq yd = 8 * 24 /9 cost = area * 14.40

OpenStudy (anonymous):

oh, i get it now. thanks so much. The answer is $307.20!